OpenStudy (anonymous):

write an indirect proof proving that if x = 40, then 2/5 x - 7 ≠ 13

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

thanks wedsite im gonna go fail now

OpenStudy (anonymous):

why wont you help me

OpenStudy (anonymous):

To do an indirect proof of A implies B, we prove something that is logically equivalent, in this case not B implies not A. For this problem,\[\frac{2}{5}x - 7 = 13 \implies x=50 \implies x \ne 40\]

OpenStudy (anonymous):

so how do i write that in a sentence form

OpenStudy (anonymous):

We are trying to show A implies B indirectly, so we show not B implies not A. In this case, not B is 2/5 x - 7 = 13. We solve this to find x = 50, which in turn tells us x is not 40, which is not A. That means A implies B, and you have proved if x = 40, then 2/5 x - 7 ≠ 13.